OpenStudy (anonymous):

A circle has a radius of 5 feet and a central angle AOB that measures 10 degrees. What is the length of the intercepted arc AB? Use 3.14 for pi and round your answer to the nearest tenth. A) 0.1 ft B) 0.3 ft C) 0.4 ft D) 0.9 ft

zepdrix (zepdrix):

What'chu think Lyshawn? :)

zepdrix (zepdrix):

Again, we're looking for a portion of the circumference \(\large\rm C=2\pi r\) 10 of the full 360 this time,\[\Large\rm S=\frac{10}{360}\cdot 2 \pi r\]

zepdrix (zepdrix):

I'm not sure what type of calculator you're using, but here is a good way to input it maybe, all at once,\[\Large\rm (10\div360)\times2\times\pi\times5\]

OpenStudy (alekos):

i thought you would have been able to work these out by now Lyshawn?

zepdrix (zepdrix):

Ya, come on girrrrrrl! You can do it! :)

OpenStudy (anonymous):

I told yall I seriously do not get math its so sad! and im still getting the wrong answer I just got 0.87......ohhhhhh or is that supposed to be rounded to 0.9 lol? lol don't judge me!

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (alekos):

well done

OpenStudy (anonymous):

lol thanks guys just bear with me please!